# Latent Diffusion Model-Enabled Real-Time Semantic Communication Considering Semantic Ambiguities and Channel Noises

### v1

@@ V-D Channel Denoising Performance
The presence of varying fading gains and noise with uncertain SNRs in wireless channels can severely impair the efficacy of SemCom systems. Accordingly, denoising the noisy signals subsequent to equalization at the receiver emerges as a vital approach to safeguard the desired meaning of the transmitted data. Typically, the channel denoising results of Deep JSCC, JPEG2000+LDPC, VE-LDM, and the proposed EECD methods are demonstrated in Fig. 9 under the conditions of both AWGN 10dB and Rayleigh 10dB. Herein, the conventional JPEG2000+LPDC approach configures CBR at 1/3 for MNIST and 0.05 for AFHQ/DIV2K datasets, whereas the CBR for DL-based methods is set at 1/16 for MNIST, and approximately 0.02 for AFHQ/DIV2K. It is noted that JPEG2000+LDPC suffers from partial bit errors and image blurring at a noise level of 10dB SNR, resulting in a lower MS-SSIM and a higher LPIPS than DL-based methods.
@@
Advancing further into DL-based methods, SemCom systems constructed on DMs and GANs outperform those based on a VAE-based Deep JSCC approach. As depicted in Fig. 9 , the Deep JSCC method exhibits a slight deficiency in certain image details relative to the latter two methods, leading to marginally inferior semantic metrics. Most crucially, the EECD method, with a subsequence length of $s$=2 used for comparison, demonstrates that the EECD methodology based on VE-LDM distillation virtually matches the performance of the original teacher model under 10dB SNR noise levels, unequivocally attesting to the effectiveness and superiority of the proposed end-to-end human perception metric-based distillation strategy.
@@
Figure 11: Semantic metrics (MS-SSIM and LPIPS) of JPEG2000+LDPC, Deep JSCC, VE-LDM, and EECD methods under different SNRs and channel states within DIV2K dataset.
@@
Numerical experiments conducted on the AFHQ dataset have provided ample validation for four distinct methodologies, revealing variations in two pivotal semantic metrics under various channel conditions and noise levels. Specifically, the CBR for JPEG2000+LDPC is set at 0.05, while a unified CBR of 0.02 is employed for the other DL-based methods. Notably, the EECD method employs different subsequence lengths of $s$=2 and $s$=1 to validate its denoising proficiency. As illustrated in Fig. 10, a perceptible degradation in quality is observed for all methods in the low-SNR area, with a particularly pronounced decline under Rayleigh and Rician channels, likely induced by fading gains. Conventionally, joint compression and error correction methods exhibit slightly inferior performance compared to DL-based approaches across varying SNRs and channel types. Furthermore, the VAE-based Deep JSCC method converges to a different perceptual quality level when compared to methods utilizing DMs and generator as SNR gradually increases. In contrast, VE-LDM and EECD methods converge to the same level of perceptual quality in high-SNR area. Most importantly, the performance of EECD can be further approximated to that of the teacher model, VE-LDM, with increased resampling lenghth, even in low-SNR area.
@@
Analogously, numerical experiments have been conducted within the DIV2K dataset, the results of which are depicted in Fig. 11. The CBR and CSI settings for four methods are consistent with those utilized in the numerical experiments for the AFHQ dataset. Specifically, the human-perceptual metrics for the denoising outcomes in the AWGN channel are superior to those in the Rayleigh and Rician channels. With regard to different channel denoising approaches, the original channel denoising DM undoubtedly achieved the most favorable performance, closely followed by the EECD curves with two different subsequence lengths, where the outcomes with $s$=2 were highly proximate to the denoising effects of the original VE-LDM, ensuring the normal transmission of semantic information.
@@
Figure 12: Semantic metrics (MS-SSIM and LPIPS) of JPEG2000+LDPC, Deep JSCC, VE-LDM, and EECD methods under different CBRs within AFHQ dataset.
@@
CBR is also an exceedingly crucial metric in SemCom, representing the ratio of transmitted bits/symbols to the size of the source data, and defining the demand for communication resources. Generally, an exemplary SemCom system is expected to maintain good reconstruction perceptual quality at a lower CBR, ultimately conserving communication bandwidth and reducing the communication burden. Fig. 12 presents the changes in average perceptual metrics for four different methods within the AFHQ dataset at CBRs ranging from 0.01 to 0.05. On one hand, the decoding quality of the conventional JPEG2000+LDPC method is heavily influenced by the compression ratio, with different CBRs potentially resulting in a manifold change in perceptual metrics. On the other hand, DL-based methods are less affected by CBR, indicating that DL-based models are robust and excel at extracting data features in the low-CBR area. Moreover, the channel denoising methods constructed based on DMs have attained optimal performance under various CBR conditions.
@@ V-E Computational Complexity Analysis
Another paramount requirement for SemCom systems is low-latency communication, encompassing minimal data processing time for encoding, transmission, denoising, and decoding. This aspect significantly impacts the age of information (AoI), quality of service (QoS), and quality of experience (QoE). The introduction of the EECD method enables the distillation of the multi-step denoising process in the latent space of the original diffusion model into a few, or even a one-step sampling process, with only a slight perceptual quality trade-off, thus facilitating real-time SemCom. Specifically, since the VE-LDM and EECD methods both utilize the same robust encoder and generator, only the computational complexity of the denoising process is analyzed. As discussed in [42], the noise prediction time complexity for the denoising U-Net used by the diffusion model is
@@
where $L$ is the number of layers, $h_{l}w_{l}$ denotes the feature size and $h_{l}w_{l}\propto H^{2}$, $C_{l}$ and $C_{l-1}$ are the number of convolutional kernels in the $l$-th and $l-1$-th layer, and $K$ is the edge length of the convolutional kernel. The channel denoising task requires only $m$ NFE based on the noise level, hence the sampling time complexity is $\textrm{Time}\sim m\times\mathcal{O}\left(\sum^{L}_{l=1}h^{2}_{l}w^{2}_{l}\cdot C_{l}\cdot C_{l-1}\cdot K^{2}_{l}\right)$. Evidently, VE-LDM may not meet the real-time denoising requirements in scenarios with low-SNR, high-CBR, or high-resolution. As demonstrated by the time consumptions for different datasets under various CSI conditions in Table III, the computing time of VE-LDM may vary dramatically according to the noise level. However, after the application of EECD, where the denoising steps are fixed at the setting value ($s$=2), the overall time required for the encoding, denoising, and decoding sequence is substantially reduced to mere tens of milliseconds.
TABLE III: Time consumption for channel denoising under different CSIs using the variance-exploding based latent diffusion model and the proposed EECD method (VE-LDM/EECD (miliseconds))
@@
Within the AFHQ dataset, the time consumption variability of VE-LDM and EECD models trained at different CBRs across a SNR range from 0dB to 20dB is illustrated in Fig. 13. In conjunction with the data presented in Table III and Fig. 13, it is evident that both the CBR and the image resolution can influence the channel denoising processing time. However, the predominant factor affecting the VE-LDM during the denoising process is the noise level. Consequently, in contrast to the VE-LDM’s more substantial and variable time complexity during denoising, the proposed EECD method consistently maintains the time required for single image transmission task within the scale of tens of milliseconds. Additionally, in line with the numerical results previously discussed, EECD does not significantly degrade semantic quality across various CSI scenarios.
@@ VI Conclusion
This paper introduces a wireless semantic communication (SemCom) system tailored to navigate the challenges of semantic ambiguities and channel noises. The proposed SemCom system’s proficiency in feature extraction diminishes the adverse effects of outliers in source data on deep learning-based communication systems and exhibits an impressive aptitude for rapid adaptation to out-of-distribution data, thereby augmenting the human-perceptual quality of decoded data. In the realm of data transmission, the advanced end-to-end consistency distillation (EECD) strategy facilitates real-time channel denoising across various pre-estimated channel state information (CSI) scenarios, achieving this with minimal perceptual quality degradation when contrasted with the existing channel denoising diffusion model techniques. Nonetheless, the real-time SemCom system based on diffusion models with unknown CSI and images with ultra-high resolution (2K/4K/6K) still warrants further investigation. Additionally, the integration of diffusion models into subsequent-generation communication paradigms, specifically goal/task-oriented SemCom systems, poses an intriguing and significant topic for future exploration.
A. Variational Upper Bound Transformation
@@
B. Training Process of VAE-WGAN-GP
The training process is illustrated in Algorithm 5.

### v4

@@ V-D Channel Denoising Performance
The presence of varying fading gains $\bm{H}_{z}$ and noise with uncertain SNRs $\bm{H}_{n}\sigma\epsilon$ in wireless channels can severely impair the efficacy of SemCom systems. Accordingly, denoising the noisy signals subsequent to equalization at the receiver emerges as a vital approach to safeguard the desired meaning of the transmitted data. Typically, the channel denoising results of Deep JSCC, JPEG2000+LDPC, VE-LDM, and the proposed EECD methods are demonstrated in Fig. 8 under the conditions of both AWGN and Rayleigh channel at SNR of 10 dB. Herein, the conventional JPEG2000+LDPC approach configures CBR at 1/3 for MNIST and 0.05 for AFHQ/DIV2K datasets for higher performances, whereas the CBR for DL-based methods is set at 1/16 for MNIST, and approximately 0.02 for AFHQ/DIV2K. It is noted that JPEG2000+LDPC suffers from partial bit errors and image blurring at a noise level of 10 dB SNR, resulting in a lower MS-SSIM and a higher LPIPS than DL-based methods.
@@
Advancing further into DL-based methods, SemCom systems constructed on DMs and GANs outperform those based on a CNN-based Deep JSCC approach. As depicted in Fig. 8, the Deep JSCC method exhibits a slight deficiency in certain image details relative to the latter two methods, leading to marginally inferior semantic metrics. Most crucially, the EECD method, with a subsequence length of $s$=2 used for comparison, demonstrates that the EECD methodology based on VE-LDM distillation virtually matches the performance of the original teacher model at SNR of 10 dB, clearly demonstrating the effectiveness and superiority of the proposed end-to-end human perception metric-based distillation strategy.
@@
Figure 10: Semantic performance metrics of JPEG2000+LDPC, Deep JSCC, DDIM, VE-LDM, and EECD methods under different SNRs and channel states within DIV2K dataset.
@@
Numerical experiments conducted on the AFHQ dataset have provided ample validation for the four distinct methodologies, revealing variations in two pivotal semantic metrics under various channel conditions and noise levels. Specifically, the CBR for JPEG2000+LDPC is set at 0.05, while a unified CBR of 0.02 is employed for the other DL-based methods, and $K$ in Rician channels is 2. Notably, the EECD method employs different subsequence lengths of $s$=2 and $s$=1 to validate its denoising proficiency. As illustrated in Fig. 9, with the SNR range of 0 dB to 20 dB, a perceptual degradation in quality is observed for all methods in the low-SNR area, with a particularly pronounced decline under Rayleigh and Rician channels, likely induced by fading gains. Evidently, all DL-enabled denoising approaches effectively address the issue of noise sensitivity present in the modulation and demodulation processes of 256-QAM, suppressing the cliff effect found in traditional communication systems while maintaining good semantic accuracy. Conventionally, joint compression and error correction methods exhibit slightly inferior performance compared to DL-based approaches across varying SNRs and channel types, even with higher CBR. Furthermore, the CNN-based Deep JSCC method converges to a different perceptual quality level when compared to methods utilizing DMs and generator as SNR gradually increases. In contrast, VE-LDM and EECD methods converge to the same level of perceptual quality in high-SNR area. Most importantly, the performance of EECD can be further approximated to that of the teacher model, i.e., VE-LDM, with increased resampling length, even in low-SNR area. The experimental results also show that the proposed end-to-end semantic metric-guided consistency training strategy significantly outperforms DDIM in low-SNR conditions with the same number of sampling steps.
@@
Similarly, experiments have been conducted within the DIV2K dataset, the results of which are depicted in Fig. 10. The CBR and CSI settings for the four methods are consistent with those utilized in the experiments for the AFHQ dataset. Overall, the semantic performance with the DIV2K dataset is slightly inferior to that with the AFHQ dataset. Among these denoising methods, the DM-enabled approaches exhibit the most robustness across different SNR levels, achieving MS-SSIM values of 13-16 dB and LPIPS values of 0.15-0.20 under 10 dB SNR conditions. Specifically, the human-perceptual metrics for the denoising outcomes with the AWGN channel are superior to those with the Rayleigh and Rician channels. In both AWGN and Rician channels, the performance of the denoising methods stabilizes at 20 dB, whereas the performance in the Rayleigh channel continues to fluctuate rapidly as the SNR increases. With regard to different channel denoising approaches, the original channel denoising DM undoubtedly achieves the most favorable performance with $m$ denoising steps, closely followed by the EECD curves with two different subsequence lengths, where the outcomes with $s$=2 are highly proximate to the denoising effects of the original VE-LDM, ensuring the normal transmission of semantic information. Additionally, EECD demonstrates the superiority of distillation and semantic learning compared to DDIM with $s=2$ in low SNR regions, while their performance is similar in high SNR regions.
@@
Figure 11: Semantic metrics of JPEG2000+LDPC, Deep JSCC, VE-LDM, and EECD methods under different CBRs within AFHQ dataset.
@@
Generally, an exemplary SemCom system is expected to maintain good reconstruction perceptual quality at a lower CBR, ultimately conserving communication bandwidth and reducing the communication burden. For different CBRs, Fig. 11 presents the changes in average perceptual metrics for the four different methods within the AFHQ dataset at CBRs ranging from 0.01 to 0.05. On one hand, the decoding quality of the conventional JPEG2000+LDPC method is heavily influenced by the compression ratio, with different CBRs potentially resulting in a manifold change in perceptual metrics. On the other hand, DL-based methods are less affected by CBR, indicating that DL-based models are robust and excel at extracting data features in the low-CBR area. Moreover, the channel denoising methods constructed based on DMs have attained superior performance under various CBR conditions.
@@ V-E Computational Complexity Analysis
Another paramount requirement for SemCom systems is low-latency communication, encompassing minimal data processing time for encoding, transmission, denoising, and decoding. The introduction of the EECD method enables the distillation of the multi-step denoising process in the latent space of the original DM into a few, or even a one-step sampling process, with only a slight perceptual quality trade-off, thus facilitating real-time SemCom. Specifically, since the VE-LDM and EECD methods both utilize the same robust encoder and generator, only the computational complexity of the denoising process is analyzed. As discussed in [38], the noise prediction computational complexity for the denoising U-Net used by the DM is
@@
where $L$ is the number of layers, $h_{l}w_{l}$ denotes the feature size and $h_{l}w_{l}\propto k$, $C_{l}$ and $C_{l-1}$ are the number of convolutional kernels in the $l$-th and $l-1$-th layer, and $K_{l}$ is the edge length of the convolutional kernel in the $l$-th layer. The channel denoising task requires only $m$ NFE based on the noise level, hence the sampling computational complexity is $\textrm{Time}\sim m\times\mathcal{O}\left(\sum^{L}_{l=1}h^{2}_{l}w^{2}_{l}\cdot C_{l}\cdot C_{l-1}\cdot K^{2}_{l}\right)$. As demonstrated by the time consumptions for different datasets under various CSI conditions in Table III, the computing time of VE-LDM may vary dramatically according to the noise level. However, after the application of EECD, where the denoising steps are fixed at the setting value ($s$=2), the overall time required for the encoding, denoising, and decoding sequence is substantially reduced to mere tens of milliseconds. In comparison with CDDM, if CBR and the number of layers of encoder, decoder and DM are the same, the reduced computational complexity of EECD is $(m-s)\cdot\mathcal{O}\left(\sum^{L}_{l=1}h^{2}_{l}w^{2}_{l}\cdot C_{l}\cdot C_{l-1}\cdot K^{2}_{l}\right)$. Inevitably, the proposed method increases the computational complexity of and the number of layers of encoder, decoder and DM are the same, the reduced computational complexity of EECD is $s\cdot\mathcal{O}\left(\sum^{L}_{l=1}h^{2}_{l}w^{2}_{l}\cdot C_{l}\cdot C_{l-1}\cdot K^{2}_{l}\right)$ compared to Deep JSCC.
TABLE III: Time consumptions of VE-LDM and EECD methods under different CSIs and datasets (VE-LDM/EECD (miliseconds))
@@
Within the AFHQ dataset, the time consumption variability of VE-LDM and EECD models trained at different CBRs across an SNR range from 0 dB to 20 dB is illustrated in Fig. 12. In conjunction with the data presented in Table III and Fig. 12, it is evident that both the CBR and the image resolution can influence the channel denoising processing time. Consequently, VE-LDM may not meet the real-time denoising requirements in scenarios with low-SNR, high-CBR, or high-resolution. However, the predominant factor affecting the VE-LDM during the denoising process is the noise level. Consequently, in contrast to the VE-LDM’s more substantial and variable time complexity during denoising, the proposed EECD method consistently maintains the time required for the denoising task within the scale of tens of milliseconds. Additionally, in line with the numerical results previously discussed, EECD does not significantly degrade semantic quality across various CSI scenarios.
In summary, the proposed SemCom system achieves a good balance between latency and robustness. The VAE (6 convolutional layers) and GAN models (7 deconvolutional layers), with average encoding and decoding times of 6.21 ms and 8.33 ms, respectively, for a single AFHQ image, introduce minimal decoding computational burden at the resource-rich transmitter. When significant semantic ambiguity is detected, the activation of one-shot learning with an average duration of approximately 450 ms ensures a rapid improvement in the quality of out-of-domain images. Fortunately, these strategies can be further enhanced by integrating advanced edge-cloud collaborative methods [63] and optimized encoding/decoding mechanisms [64].
@@ VI Conclusion
This paper introduces a wireless semantic communication (SemCom) system tailored to navigate the challenges of semantic ambiguities and channel noises. The proposed SemCom system’s proficiency in feature extraction diminishes the adverse effects of outliers in source data on deep learning-based communication systems and exhibits an impressive aptitude for rapid adaptation to data with unknown distribution, thereby augmenting the human-perceptual quality of decoded data. In the realm of data transmission, the advanced end-to-end consistency distillation (EECD) strategy facilitates real-time channel denoising across various pre-estimated channel state information (CSI) scenarios, achieving this with minimal perceptual quality degradation when contrasted with the existing channel denoising diffusion model techniques. Nonetheless, the real-time SemCom system based on diffusion models with unknown CSI, images with ultra-high resolution (2K/4K/6K), and large network environments still warrants further investigation. Additionally, the integration of diffusion models into next-generation communication paradigms, specifically goal/task-oriented SemCom systems, poses an intriguing and significant topic for future exploration.
-A VUB Transformation for VAE-WGAN
@@
-B Training Process of VAE-WGAN-GP
The training process of deep CNN based VAE-WGAN with gradient penalty [65] is illustrated in Algorithm 5, where $\alpha_{\bm{\phi}}$ and $\alpha_{\bm{\psi}}$ are the loss balance hyperparameters.
@@ 9      Sample $\bm{x}\sim q(\bm{x})$ and $\bm{z}\sim q_{\bm{\psi}}(\bm{z})$;
-C Proof of Robust Encoder’s VUB
According to conditional Markov random field model [66], the joint distribution of data $\bm{x}$ and data with semantic error $\bm{x}+\bm{\delta}$ is
$p_{\bm{\psi}}(\bm{x},\bm{x}+\bm{\delta})\propto$ (33)
[…]
where $\beta$ denotes the nonnegative coupling parameter. Consequently, considering the joint distribution $q(\bm{z},\bm{z}^{\prime\prime})$, the evidence lower bound has the following form
$\mathbb{E}_{q}\left[\log p_{\bm{\psi}}(\bm{x},\bm{x}+\bm{\delta})\right]\geq\mathbb{E}_{q(\bm{z})}\left[\log p_{\bm{\psi}}(\bm{x}|\bm{z})\right]+\mathbb{E}_{q(\bm{z})}\left[\log p_{\bm{\psi}}(\bm{z})\right]$ (34)
$+\mathbb{E}_{q(\bm{z}^{\prime\prime})}\left[\log p_{\bm{\psi}}(\bm{x}+\bm{\delta}|\bm{z}^{\prime\prime})\right]+\mathbb{E}_{q(\bm{z}^{\prime\prime})}\left[\log p_{\bm{\psi}}(\bm{z}^{\prime\prime})\right]$
$-\frac{\beta}{2}\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\bm{d}(\bm{z},\bm{z}^{\prime\prime})+\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\left[\log q(\bm{z},\bm{z}^{\prime\prime})\right].$
To decode clean data without changing the decoder parameters $\bm{\psi}$, by simply integrating out $\bm{x}+\bm{\delta}$ in Eq. (34) and according to Jensen’s inequality, term $\mathbb{E}_{q}\left[-\log p_{\bm{\psi}}(\bm{x})\right]$ can be transformed into Eq. (10).